OpenStudy (anonymous):

(x-3)^2/9 + (y+1)^2/4. the center is (3,-1).....so why is the foci not (3,-1+sqrt5),(3,-1-sqrt5)? my answer for foci is wrong and i don't know why.

OpenStudy (anonymous):

i think your orientation is wrong

OpenStudy (anonymous):



OpenStudy (anonymous):

so instead of what you had, it is \(\sqrt{5}\) units left and right, not up and down

OpenStudy (anonymous):

hope that is clear, and you know it because \(9>4\)

OpenStudy (anonymous):

ie (3,-1+sqrt5) should be (-1, 3-sqrt5)

OpenStudy (anonymous):

i hate to disagree with the great estudier, but i am afraid since it is to the left and right, it should be \[(3-\sqrt{5},-1), (3+\sqrt{5},-1)\]

OpenStudy (anonymous):

Oops:-) (I did give u the medal)

OpenStudy (anonymous):

merci (that is how i knew to look!)

OpenStudy (anonymous):

thank you both!!